OpenStudy (anonymous):

Use the co-factor expansion of determinants to evaluate the determinants of the following n x n matrices. (matrices in comments) Evaluate the determinants with n = 14 and input them as your answers.

OpenStudy (anonymous):

\[A = (a _{ij}) = \left[\begin{matrix}0 & 0 & ... & 0 & 1 \\0 & 0 & ... & 2 & 0 \\ ... & ... & ... & ... & ... \\0 & n-1 & 0 & ... & 0 \\n & 0 & 0 & ... & 0 \\\end{matrix}\right]\]\[B = (b _{ij}) = \left[\begin{matrix}0 & 1 & 0 & ... & 0 \\0 & 0 & 2 & ... & 0 \\... & ... & ... & ... & ... \\0 & 0 & 0 & ... & n-1 \\n & 0 & 0 & ... & 0 \\\end{matrix}\right]\]

OpenStudy (anonymous):

if n = 14, then det(A) = ? if n = 14, then det(B) = ?